OpenStudy (anonymous):

A plane flies 465 miles with the wind and 345 miles against the wind in the same length of time. If the speed of the wind is 20 mph, find the speed of the plane in still air

OpenStudy (anonymous):

t(a+w)=465 t(a-w)=345 w=20 t(a+20)=465 t(a-20)=345 ta +20t =465 ta -20t =345

OpenStudy (anonymous):

When solved, I get t=3, a=135

OpenStudy (anonymous):

a= speed of plane relative to still air